OpenStudy (anonymous):

How do you do this? Put in increasing order. 7/10, 1/2, 3/4?

OpenStudy (anonymous):

You must find the least common multiple wich is 40 (7/10)*4 = 28/40 (1/2)*20 = 20/40 (3/4)*10 = 30/40 now its easy